OpenStudy (anonymous):

Braiden wants to add a rectangular concrete slab to existing drivway. How much concrete should he order to make a 14 feet long, 8 1/2 feet wide and 8 inches thick slab?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here, we have to compute the volume of the concrete slab. Now, the volume V of a parallelepiped whose side are a, b, and c is given by the subsequent formula: \[\large V = a \times b \times c\] In your case we have: \[\large \begin{gathered} a = 14 \hfill \\ b = 8\frac{1}{2} \hfill \\ c = 8 \hfill \\ \end{gathered} \] so, what is V?

OpenStudy (anonymous):

952

OpenStudy (michele_laino):

oops...I have made an error, we have to exprees all of the sides in inches, so: \[\large \begin{gathered} a = 14 \times 12 = ...? \hfill \\ b = \left( {8\frac{1}{2}} \right) \times 12 = ... \hfill \\ c = 8 \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

express*

OpenStudy (anonymous):

a=168 b= 102 c= 8

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

so: \[\large V = a \times b \times c = 168 \times 102 \times 8 = ...?\]

OpenStudy (anonymous):

137,088

OpenStudy (michele_laino):

yes! please remember its unit of measure is inches^3

OpenStudy (anonymous):

okay:) so whats the next step")

OpenStudy (michele_laino):

we have finished!

OpenStudy (anonymous):

The answer book says 79.33

OpenStudy (michele_laino):

since the volume is measured in feet^3 so in order to get your answer, we have to divide 137,088 by 12^3= 1728, then we have: \[\large V = a \times b \times c = \frac{{168 \times 102 \times 8}}{{1728}} = ...?\]

OpenStudy (anonymous):

great!!! 79.33:)

OpenStudy (michele_laino):

:)